# Characterization of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

### v1

@@ III.2 Constructing birefringence maps from TWE
The biggest challenge for birefringence characterizations of KAGRA’s current ITMs comes from the fact that previously TWE measurement setup was not optimized for birefringence study. According to Equation 20 in Section II.3, in order to extract birefringence information, we need subtractions of two pairs of TWE maps and then take the division. In typical Fizeau interferometer measurements, low order effects like the piston and tilt of the surface are not measured. When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample. This means that the subtraction of two maps will contain unknown piston and tilt. The numerator and denominator in Equation 20 containing these unknown terms will give wrong birefringence information.
@@
The center of the map also changes after rotating the sample. But we found the decentering is very small (less than 2 mm) and its effect can be negligible when compared with effects from piston and tilt. In addition, we found astigmatisms in original TWE maps. The left plot in Figure 4 shows a TWE map after removing the curvature. A clear astigmatism term can be seen from the map, which is plotted in the right. We found the rotation of this astigmatism does not change when rotating the sample. So, we think the astigmatism comes from the gravity induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer. The astigmatism is removed for each TWE map individually.
@@
Figure 4: Raw TWE map containing astigmatism which mainly comes from deformation due to gravity. As it is not exactly in the vertical direction, we think the surface of the reference sphere used in the Fizeau interferometer is slightly ellipsoidal. After removing the astigmatism, the TWE is shown in the top left plot in Figure 3.
@@
From Equation 21, $\theta$ is wrapped in the range of ($-45^{\circ}$, $45^{\circ}$) due to $\arctan$ function and we need to unwrap the data. Phase unwrapping techniques are widely used in optical metrology and interferometry [32, 33]. In our case, the problem of unknown piston and tilt is that it will give wrong $\theta$ and cause failure of unwrapping. There are algorithms improving phase unwrapping robustness again noise [34, 35, 36, 37]. Most of these methods require the noise to be small compared with the phase change. But the unmeasured piston and tilt in our TWE map can be large. As we take the division of two subtractions, the error will be further amplified when the denominator in Equation 20 is close to zero. The left plot in Figure 5 shows our initially derived $\theta$. The unwrapping algorithm cannot work properly for such kind of map (shown in the right). This results in that the unwrapped surface is discontinuous. The algorithm used in this paper is according to [38, 39].
@@
Figure 5: The left plot shows initially derived $\theta$ map according to Equation 20. Unknown piston and tilt will give wrong $\theta$ and cause failure of unwrapping shown in the right where the unwrapped surface is discontinuous. When performing parameter search for pistons, the 2D matrix is expended to 1D vector along the arrow direction. Each arrow represents a row in the 477$\times$477 data matrix.
@@
In order to recover the correct $\theta$, we perform parameter search for piston terms in the numerator and denominator in Equation 20 (while tilt terms are removed individually for each TWE map and ignored here). This yields
@@
where $p_{1}$ and $p_{2}$ are the missing piston terms. For each pair of $p_{1}$ and $p_{2}$ values, we unwrap the derived raw $\theta$ map. Here, we assume the unmeasured pistons are the main cause for the failure of unwrapping. The two-dimensional unwrapped map is then expended to one-dimensional data along the arrow direction shown in the right in Figure 5. If the unwrapped map is a continuous surface, the one-dimensional data should be a continuous curve. So the overall continuity factor of the map can be simply described by
@@
where $\epsilon_{k}$ is the TWE value at each point of the one-dimensional data and $n$ is the total number of the data. Lower value of $\mathcal{C}$ means a smoother surface while higher value of $\mathcal{C}$ means more value jumps over the surface (see the right plot in Figure 5). Continuity factors of ITMX and ITMY are shown in the right in Figure 6. For each pair of $p_{1}$ and $p_{2}$, $\theta$ and $\alpha_{-}$ can be derived. Then we calculate the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ in Equation 2. This gives the field of linearly polarized light in perpendicular direction with the input polarization. It’s beam shape is shown in the left in Figure 6 with different piston values.
Figure 6: Parameter scan for piston term for ITMX (top) and ITMY (bottom). Plots in the left are single-bounce reflected (passing through the mirror substrate twice) beam shapes in p-polarization (the input is pure s-polarization) with different pistons applied to subtractions of TWEs. Numbers in white are the ratio percentage of beam power in p-polarization over the total input power. Plots in the right are continuity factors of $\theta$ map after unwrapping. It can be seen that $p_{1}$ mainly influence the continuity factor while $p_{2}$ determines the beam shape and power. We search for the correct pistons by eye where continuity factors are small and beam shapes match with measurements shown in Figure 8. We think the beam shape shown in red square gives the best estimation as a consideration of continuity factor and measured beam shape. But actually beams in the green area all have similar shapes with power varying a lot.
@@ III.3 Comparison with in-situ measurements
We have performed birefringence measurements for the sapphire ITMs in KAGRA. A simplified schematic of KAGRA configuration is shown in Figure 7. The laser after the Faraday Isolator is in pure s-polarization. The Power Recycling Mirror (PRM), Signal Recycling Mirror (SRM) and End Test Masses (ETMX and ETMY) are all misaligned so that the beam is reflected back only by ITMs. The plot shows the birefringence measurement for ITMX where ITMY is misaligned. Coatings of the beam splitter (BS) are optimized for s-polarization with a 50%:50% ratio for reflection and transmission at ${}^{\circ}$ incidence. The ratio for p-polarization is 20%:80% according to the coating design. Since there are no direct measurements for p-polarization, practical reflection and transmission may be different from design values. The reflected beams from ITMs are measured at POP and POS ports. There are photodetectors (PD) and cameras for both s-polarization and p-polarization at POP, while there are only photodetectors at POS.
@@
Table 1: Comparison between the measured beam power in p-polarization at POP and POS ports with calculations from $\theta$ and $\alpha_{-}$ maps. Numbers are percentages of the p-polarization power among total power reflected from ITMX or ITMY.
@@
The measured beam power in p-polarization (as a percentage of the total reflected power) is shown in Table 1, where the effect of unbalanced reflection and transmission of the BS for p-polarization is calibrated (according to coating design values). In principle, the measured beam power content in p-polarization should be the same for POP and POS ports. But in our measurements they are different from each other. We think one reason could be that the practical reflection and transmission of the BS for p-polarization deviate from design. Another reason could be due to the clipping and ghost beam. When the measurement was done, gate valves were added to vacuum tubes whose diameters are just slightly bigger than the beam size. We also found there were a lot of ghost and scattering beams at POP and POS. Though we dumped most of them, there may be still some left and seen by the photodetectors.
@@
Measured beam shapes are shown in the right in Figure 8. By comparing our calculated beam shapes from the previous section with practical measurements, we can find the correct pair of $p_{1}$ and $p_{2}$ by eye which gives both low continuity factor and good agreement with measured beam shapes (see Figure 6). The beam power in p-polarization are then calculated (values are shown in Table 1) with the most suitable pair of $p_{1}$ and $p_{2}$ (marked with red square in Figure 6). The estimated beam power from our maps is smaller than measurements. We think one of the reason is that the unknown tilt term in the measured TWE maps will influence the beam power. By adding tilt influences in the parameter scan, it is possible to increase the calculated power in p-polarization while keeping the beam shape unchanged (or changed by little). In this paper, we don’t perform the full parameter scan for tilt, since the purpose of this paper is to show the feasibility of the method of extracting birefringence information from TWE measurements. An example of the influence of tilt is presented in the Appendix LABEL:app:_tilt.
@@
Figure 8: Comparison between calculated p-polarization beam shape and KAGRA measurements. Dashed circles in white represent the Gaussian beam diameter.
@@
Another factor that can affect our estimated power is the beam spot position on the mirror. The beam spots on ITMs in KAGRA are not exactly in the center but a couple of centimeters away where the power recycling gain and arm transmission power are higher, which is probably due to the homogeneous birefringence effect. Figure 9 shows our estimation how much the beam power in p-polarization will change when the beam is miscentered.
@@
Figure 9: Power of p-polarization when the beam spot is moved away from the mirror center. The x-axis represents the direction of the decenter and $r$ is the distance between the beam spot center and mirror center. Birefringence effect gets smaller when moving the beam towards southeast on ITMX and towards north on ITMY, which can also be seen from Figure 8.
@@
The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10. From these maps, we can see the birefringence in KAGRA’s current ITMs are quite inhomogeneous. If we compare these maps with the beam shape in p-polarization shown in Figure 8, it can be found the beam shape is mainly determined by the shape of $\theta$ map. This is because $\alpha_{-}$ is small in our case. The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarization from Equation 7.
@@
Figure 10: Calibrated $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY. Dashed circles represent the laser beam diameter on mirrors. The plots shown here are the one-way differential phase $\alpha_{-}/2$. The unit of $\alpha_{-}$ is nanometer and the laser wavelength 1064 nm corresponds to a radian of $2\pi$.
@@
As discussed above, a good knowledge of piston and tilt in TWE measurements, as well as the location of the beam spot are necessary in order to precisely estimate the birefringence effect. For more accurate characterizations, the previous setup requires improvements and optimizations for birefringence study.
@@ IV Conclusion and discussion
We have established a method to reconstruct the 2D birefringence information of current KAGRA’s sapphire ITMs from TWE measurements. By combining several TWE measurements with a Fizeau interferometer using linearly polarization in different rotation angles, we are able to successfully construct the birefringence map over a 200 mm area in diameter. Our estimated beam shape for single bounce reflected field in orthogonal polarization shows a good match with in-situ measurements in KAGRA.
It is worth to point out that these TWE measurements were not intended for birefringence characterizations. More accurate reconstruction requires improvements of the previous setup which should target for birefringence study. One possible improvement could be enabling the control of the polarization rotation angle of the input laser in the Fizeau interferometer rather than rotating the tested mirror, as this will remove the influence of the low order (piston, tilt and astigmatism caused by deformation due to gravity) error term in TWE. For future gravitational wave detectors, since most potential materials feasible for cryogenic operation are birefringent, birefringence has become an important parameter for their core optics. This technique of birefringence mapping described in this paper is fast and simple. It can be crucial for future detectors where larger test masses are used. In addition, specifications of mirror birefringence requires a full understanding of how the birefringence affects the performance of the interferometer. The derived birefringence map in this paper can be used to study the effects of inhomogeneous birefringence in future detectors and help us to gain insights of this knowledge.
@@ Acknowledgements.
We would like to thank Miyoki Shinji, Akutsu Tomotada, Takafumi Ushiba and Masahide Tamaki in KAGRA for their support and help for the measurements. We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions. This work was supported by JSPS KAKENHI Grant No. JP20H05854, No. 23H01205, and by JST PRESTO Grant No. JPMJPR200B. KAGRA is supported by MEXT, JSPS Leading-edge Research Infrastructure Program, JSPS Grant-in-Aid for Specially Promoted Research 26000005, JSPS Grant-inAid for Scientific Research on Innovative Areas 2905: JP17H06358, JP17H06361 and JP17H06364, JSPS Core-to-Core Program A. Advanced Research Networks, JSPS Grantin-Aid for Scientific Research (S) 17H06133 and 20H05639 , JSPS Grant-in-Aid for Transformative Research Areas (A) 20A203: JP20H05854, the joint research program of the Institute for Cosmic Ray Research, University of Tokyo, National Research Foundation (NRF), Computing Infrastructure Project of Global Science experimental Data hub Center (GSDC) at KISTI, Korea Astronomy and Space Science Institute (KASI), and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS), AS Grid Center (ASGC) and the National Science and Technology Council (NSTC) in Taiwan under grants including the Rising Star Program and Science Vanguard Research Program, Advanced Technology Center (ATC) of NAOJ, and Mechanical Engineering Center of KEK.

### v2

@@ III.2 Constructing birefringence maps from TWE
The biggest challenge for the birefringence characterisation of KAGRA’s current ITMs is that the previous TWE measurement setup was not optimised for birefringence studies. According to Equation (20) in Section II.3, to extract the birefringence information, we need to subtract two pairs of TWE maps and then divide. In typical Fizeau interferometer measurements, low-order effects such as the piston and surface tilt are removed from the results. When doing measurements with different $\theta$, the distance and alignment between the sample and reference mirror have changed after rotating the sample. This implies that the difference between the two maps contains an unknown piston and tilt. The numerator and denominator in Equation (20), which contain these unknown terms, provide incorrect birefringence information.
@@
The centre of the map also changed after the sample was rotated. However, we observed that the decentring was very small (less than 2 mm), and its effect was negligible compared with the effects of the piston and tilt. In addition, astigmatism was observed on the original TWE maps. The left plot in Figure 4 shows the TWE map after removing the curvature. A clear astigmatism term can be observed on the map plotted on the right. We observed that the rotation of this astigmatism did not change when the samples were rotated. Therefore, we believe that the astigmatism originates from the gravity-induced deformation of the sample and the astigmatism of the reference mirror used for the Fizeau interferometer. Astigmatism was individually removed from each TWE map.
@@
Figure 4: Raw TWE map containing astigmatism which mainly occurs due to deformation due to gravity. As it is not in the vertical direction, we assume that the surface of the reference sphere used in the Fizeau interferometer is slightly ellipsoidal. After removing the astigmatism, the TWE is shown in the top left plot in Figure 3. The algorithm used in processing the maps was based on Simtools [32].
@@
From Equation (21), $\theta$ is wrapped in the range of ($-45^{\circ}$, $45^{\circ}$) due to the $\arctan$ function and we need to unwrap the data. Phase unwrapping techniques are widely used in optical metrology and interferometry [33, 34]. In our case, the problem of the unknown piston and tilt is that it will give the wrong $\theta$ and cause failure of unwrapping. Some algorithms improve phase unwrapping robustness against noise [35, 36, 37, 38]. Most of these methods require the noise to be small compared with the phase change. However, the unmeasured piston and tilt on the TWE map can be large. As we take the division of two subtractions, the error will be further amplified when the denominator in Equation (20) is close to zero. The left plot in Figure 5 shows our initially derived $\theta$. The unwrapping algorithm does not function properly for such maps (shown on the right). This resulted in a discontinuous, unwrapped surface. The algorithm used in this study was based on [39, 40].
@@
Figure 5: The left plot shows initially derived $\theta$ map according to Equation (20). The unknown piston and tilt will give wrong $\theta$ and cause failure of unwrapping shown on the right where the unwrapped surface is discontinuous. When performing parameter search for pistons, the 2D matrix is expanded to 1D vector along the arrow direction. Each arrow represents a row in the 477$\times$477 data matrix.
@@
To recover the correct $\theta$, we perform parameter search for piston terms in the numerator and denominator shown in Equation (20) Tilt terms are removed individually for each TWE map (by removing first-order Zernike polynomials) and ignored here. This yields
@@
where $p_{1}$ and $p_{2}$ are missing piston terms. For each pair of $p_{1}$ and $p_{2}$ values, the derived raw $\theta$ map was unwrapped. Here, we assume that unmeasured pistons are the main cause of unwrapping failure. The two-dimensional unwrapped map is then expanded to one-dimensional data along the arrow direction, as shown on the right in Figure 5. If the unwrapped map is a continuous surface, the one-dimensional data should be a continuous curve. Therefore, the overall continuity factor of the map can be simply described by
@@
where $\epsilon_{k}$ is the TWE value at each point of the one-dimensional data and $n$ is the total number of the data. A lower value of $\mathcal{C}$ indicates a smoother surface, whereas a higher value of $\mathcal{C}$ means that more values jump over the surface (refer the right plot shown in Figure 5). The continuity factors of ITMX and ITMY are shown on the right side of Figure 6. For each pair of $p_{1}$ and $p_{2}$, $\theta$ and $\alpha_{-}$ can be derived. Subsequently, we calculate the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ in Equation (2). This provided a field of linearly polarised light in the direction perpendicular to the input polarisation. The beam shape is shown on the left in Figure 6 with different piston values.
Figure 6: Parameter scan for piston term for ITMX (top) and ITMY (bottom). Plots on the left are single-bounce reflected (passing through the mirror substrate twice) beam shapes in p-polarisation (the input is pure s-polarisation) with different pistons applied to subtractions of TWEs. Numbers in white are the ratio percentage of beam power in p-polarisation over the total input power. Plots on the right are continuity factors of $\theta$ map after unwrapping. It can be observed that $p_{1}$ mainly influences the continuity factor, whereas $p_{2}$ determines the beam shape and power. We search for the correct pistons by eye where continuity factors are small and beam shapes match with measurements shown in Figure 8. We assume that the beam shape shown in the red square gives the best estimation as a consideration of the continuity factor and measured beam shape. However, beams in the green area all have similar shapes with significantly varying power.
@@ III.3 Comparison with in-situ measurements
Birefringence measurements were performed for the sapphire ITMs at KAGRA. A simplified schematic of the KAGRA configuration used during this measurement is shown in Figure 7. The laser after the Faraday Isolator was in pure s-polarisation. The power recycling mirror (PRM), signal recycling mirror (SRM), and end test masses (ETMX and ETMY) are all misaligned such that the beam is reflected only by the ITMs. The figure shows the particular setup for measuring the birefringence of ITMX, where ITMY is misaligned. The coatings of the beam splitter (BS) were optimised for s-polarisation with a 50%:50% ratio for reflection and transmission at 45 incidence. The p-polarisation ratio was 20%:80% according to the coating design. Because there is no direct measurement of coating reflectivity for p-polarisation, the practical reflection and transmission may be different from the design values. The beams reflected from the ITM were measured at the POP and POS ports. Photodetectors (PD) and cameras are used for both s-polarisation and p-polarisation at the POP, whereas only photodetectors are used at the POS.
@@
Table 2: Comparison between the measured beam power in p-polarisation at POP and POS ports with calculations from $\theta$ and $\alpha_{-}$ maps. Numbers are percentages of the p-polarisation power with respect to the total power reflected from the ITMX or ITMY.
@@
The measured beam power in p-polarisation (as a percentage of the total reflected power) is listed in Table 2, where the effect of the unbalanced reflection and transmission of the BS on the p-polarisation is calibrated (according to the coating design values). In principle, the measured beam power content in p-polarisation should be the same for the POP and POS ports. However, our measurements are different from each other. One reason for this could be that the practical reflection and transmission of the BS for p-polarisation deviate from the design. Another reason could be clipping and ghost beams. When the measurement was performed, gate valves with diameters slightly larger than the beam diameter were added to the vacuum tubes. We also found many ghost and scattering beams at the POP and POS. Although we dumped most of them, there may be still some remaining and are seen by the photodetectors.
@@
The measured beam shapes are shown on the right side in Figure 8. By comparing our calculated beam shapes from the previous section with practical measurements, we can find the correct pair of $p_{1}$ and $p_{2}$ visually, which gives both a low continuity factor and good agreement with the measured beam shapes (refer Figure 6). The beam power in p-polarisation was then calculated (the values are listed in Table 2) using the most suitable pair of $p_{1}$ and $p_{2}$ (marked with the red square shown in Figure 6). The beam power estimated from our maps was lower than that from the measurements. One reason is that the unknown tilt term in the measured TWE maps influences the beam power. By adding tilt influences in the parameter scan, it is possible to increase the calculated power in p-polarisation while maintaining the beam shape unchanged (or slightly changed). In this study, we did not perform a full parameter scan for tilt because the purpose was to demonstrate the feasibility of the method of extracting birefringence information from TWE measurements. An example of the influence of the tilt is presented in Appendix B.
@@
Figure 8: Comparison between calculated p-polarisation beam shape and KAGRA measurements. Dashed circles in white represent the Gaussian beam diameter, the width at which the amplitude is $1/e$ of its value in the centre.
@@
Another factor that can affect the estimated power is the position of the beam spot on the mirror. The beam spots on the ITMs in KAGRA are not exactly in the centre but are a couple of centimetres away where the power recycling gain and arm transmission power are higher, which is probably due to the inhomogeneous birefringence effect. Figure 9 shows our estimate of the extent to which the p-polarised beam power changes when the beam is miscentered.
@@
Figure 9: Power of p-polarisation when the beam spot is moved away from the mirror centre. The x-axis represents the direction of the decentre and r is the distance between the beam spot centre and mirror centre. Birefringence effect gets smaller when moving the beam towards southeast on ITMX and towards north on ITMY, as shown in Figure 8.
@@
The derived $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY are shown in Figure 10. From these maps, we can see that the birefringence in KAGRA’s current ITMs is inhomogeneous. If we compare these maps with the beam shape in p-polarisation shown in Figure 8, it can be found that the beam shape is mainly determined by the shape of $\theta$ map. This is because $\alpha_{-}$ is small in our case. The $\sin 2\theta$ term plays a more important part than $\sin\alpha_{-}$ in the amplitude of the field in p-polarisation from Equation (7).
@@
Figure 10: Calibrated $\theta$ and $\alpha_{-}$ maps for ITMX and ITMY. Dashed circles represent the laser beam diameter on mirrors. The plots shown here are the one-way differential phase $\alpha_{-}/2$. The unit of $\alpha_{-}$ is nanometre and the laser wavelength 1064 nm corresponds to a radian of $2\pi$.
@@
As discussed above, a good knowledge of the piston and tilt in TWE measurements and the location of the beam spot is necessary to precisely estimate the birefringence effect. For more accurate characterisation, the Fizeau interferometer requires improvements and optimisations for the birefringence study.
@@ IV Conclusion and discussion
We established a method for reconstructing the 2D birefringence information of current KAGRA sapphire ITMs from TWE measurements. By combining several TWE measurements with a Fizeau interferometer using linear polarisation at different rotation angles, we successfully constructed a birefringence map over a 200 mm area in diameter. Our estimated beam shape for a single-bounce reflected field in orthogonal polarisation shows a good match with the in situ measurements in KAGRA.
Note that these TWE measurements were not intended for birefringence characterisations. More accurate reconstruction requires improvements to the previous setup, which should target birefringence studies. One possible improvement could be enabling control of the polarisation rotation angle of the input laser in the Fizeau interferometer rather than rotating the tested mirror, as this will remove the influence of the low-order (piston, tilt, and astigmatism caused by deformation due to gravity) error term in the TWE. For future gravitational wave detectors, birefringence has become an important parameter for their core optics because most potential materials that are suitable for cryogenic operation are birefringent. The birefringence-mapping technique described in this study is fast and simple. This is crucial for future detectors with larger test masses. In addition, the specifications of mirror birefringence require a complete understanding of how birefringence affects the interferometer performance. The derived birefringence map in this study can be used to study the effects of inhomogeneous birefringence in future detectors and help us gain insights into this knowledge.
@@ Acknowledgements.
We would like to thank Shinji Miyoki and Tomotada Akutsu at KAGRA for their support for the measurements. We also thank Masaki Ando, Kentaro Somiya, Kentaro Komori and Satoru Takano for insightful discussions. This work was supported by JSPS KAKENHI Grant No. JP20H05854, No. 22H01228, No. 23H01205, and by JST PRESTO Grant No. JPMJPR200B. KAGRA is supported by MEXT, JSPS Leading-edge Research Infrastructure Program, JSPS Grant-in-Aid for Specially Promoted Research 26000005, JSPS Grant-inAid for Scientific Research on Innovative Areas 2905: JP17H06358, JP17H06361 and JP17H06364, JSPS Core-to-Core Program A. Advanced Research Networks, JSPS Grantin-Aid for Scientific Research (S) 17H06133 and 20H05639 , JSPS Grant-in-Aid for Transformative Research Areas (A) 20A203: JP20H05854, the joint research program of the Institute for Cosmic Ray Research, University of Tokyo, National Research Foundation (NRF), Computing Infrastructure Project of Global Science experimental Data hub Center (GSDC) at KISTI, Korea Astronomy and Space Science Institute (KASI), and Ministry of Science and ICT (MSIT) in Korea, Academia Sinica (AS), AS Grid Center (ASGC) and the National Science and Technology Council (NSTC) in Taiwan under grants including the Rising Star Program and Science Vanguard Research Program, Advanced Technology Center (ATC) of NAOJ, and Mechanical Engineering Center of KEK.